OpenStudy (rainbow_dash):

Find the first five terms of the given sequence. a_1 = 4, a_n+1 = 3a_n+12

OpenStudy (anonymous):

\[a_2=3(4)+12\]

OpenStudy (anonymous):

\[a_3=3(a_2)+12\] \[a_4=3(a)_3+12\] \[a_5=3(a_4)+12\]

OpenStudy (rainbow_dash):

ohhhh ok thanks :)